OpenStudy (anonymous):

What is the mass of 2.56 × 10–4 moles of Fe2O3? A. 4.09 × 10–2 g B. 159.6 g C. 6.23 × 105 D. 1.60 × 0–6 g

OpenStudy (anonymous):

first find the molecular weight of Fe2O3

OpenStudy (anonymous):

molecular weight of Fe2O3 =56*2 +16*3 =112+48=160

OpenStudy (anonymous):

now by definition of mole 1 mole of Fe2O3 weighs 160 g

OpenStudy (anonymous):

hence 2.56 × 10–4 moles of Fe2O3 must weigh 2.56 × 10–4 * 160 g....

OpenStudy (ash2326):

@Joshyface I think this should be posted in Chemistry. Please post chemistry question in Chemistry subject now onwards. http://openstudy.com/study#/groups/Chemistry

OpenStudy (anonymous):

@@Joshyface its 2.56*160/10000 g = 4.09 × 10–2 g so option A